# CDI: Copyrighted Data Identification in Diffusion Models

### v2

@@ Appendix G Additional Features
Appendix H Experimental Setup
H.1 Models
@@
H.2 Compute Resources and Feature Extraction Time
@@
Appendix I MIAs and their Model Access Types
@@
Table 6: Each feature with its corresponding Model Access Type.
@@
Appendix J Additional Experiments
Gray-Box Model 8000 3000 2000 4000 2000 400 200 70
White-Box Model 6000 3000 2000 2000 1000 300 200 70
J.1 Number of Model Training Steps and CDI Eeffectiveness
[…]
J.2 Model Details and CDI Effectiveness
Appendix K More Results for the Non-members in P and False Positives
Appendix L Detailed Gray-box vs. White-box comparison
@@
Appendix M Analysis of the scoring function
@@
Appendix N From p-value to TPR@FPR=1%.
@@
Appendix P MIAs Fail on DI Task
In this experiment, we compare CDI to MIAs on the task of DI. Similarly to set-level MIA we introduce in Sec. 5.2, we follow this procedure: We apply a MIA to each sample in the suspect set, returning Positive prediction if the score for any sample exceeds a certain threshold. Repeating this process for every suspect set and varying the threshold yields a ROC curve for each MIA. Finally, we compare these ROC curves against the ROC curve for CDI, obtained as described in App. O.
@@
Appendix Q Further Applicability of Our Novel Features
One of our contributions is the new features we introduce in Sec. 4.1. As we use these features to fit our scoring function, we can also use them to perform the default threshold MIA. We introduce the detailed results of that experiment in App. R. In this paragraph we analyze the characteristics of our features.
Q.1 Gradient Masking
@@
Q.2 Noise Optim and Multiple Loss
@@
Q.3 Effect of aggregation on MIA and CDI performance
@@
Appendix R Further Evaluation of MIAs

### v3

@@ H.3 Extending CDI to other data modalities
Appendix I Experimental Setup
I.1 Models
@@
I.2 Compute Resources and Feature Extraction Time
@@
Appendix J Model Details and CDI Effectiveness
Appendix K Number of Model Training Steps and CDI Effectiveness
[…]
Appendix L More Results for the Non-members in P and False Positives
[…]
$|\textbf{U}|/|\textbf{P}|$ ratio LDM256 U-ViT256 DiT256 U-ViT512 DiT512 U-ViT256-T2I U-ViT256-T2I-Deep
1.00 6000 3000 2000 2000 1000 200 70
0.90 6000 3000 2000 2000 2000 200 70
0.75 6000 3000 2000 3000 2000 200 70
0.50 6000 3000 2000 3000 2000 300 90
0.25 7000 4000 2000 4000 2000 400 200
0.10 8000 5000 4000 6000 5000 800 300
Gray-Box Model 8000 3000 2000 4000 2000 400 200 70
White-Box Model 6000 3000 2000 2000 1000 300 200 70
[…]
In practice, the training-to-test set ratio ($|U|/|P|$) is often much smaller than one. We analyze CDI’s performance in such scenarios and find it remains effective even when the victim has access to fewer test samples ($|U|<|P|$). We report the number of samples in $P$ needed to reject $H_{0}$ for given ratios ($|U|=\text{ratio}\times|P|$). Notably, an extreme set imbalance can be mitigated by increasing $P$.
Appendix N MIAs and their Model Access Types
@@
Table 11: Each feature with its Model Access Type.
@@
Appendix O Gray-box vs. White-box comparison
@@
Appendix P Analysis of the scoring function
@@
Appendix Q From p-value to TPR@FPR=1%.
@@
Appendix S MIAs Fail on DI Task
In this experiment, we compare CDI to MIAs on the task of DI. Similarly to set-level MIA we introduce in Sec. 5.2, we follow this procedure: We apply a MIA to each sample in the suspect set, returning Positive prediction if the score for any sample exceeds a certain threshold. Repeating this process for every suspect set and varying the threshold yields a ROC curve for each MIA. Finally, we compare these ROC curves against the ROC curve for CDI, obtained as in App. R.
@@
Appendix T Further Applicability of Our Novel Features
One of our contributions is the new features introduced in Sec. 4.1. As we use these features to fit our scoring function, we can also use them to perform the default threshold MIA. We introduce the detailed results of that experiment in App. U. Here, we analyze the characteristics of our features.
T.1 Gradient Masking
@@
T.2 Noise Optim and Multiple Loss
@@
T.3 Effect of aggregation on MIA and CDI performance
@@
Appendix U Further Evaluation of MIAs